OpenStudy (anonymous):

If 3a + 4b = 4a - 4b = 21, find the value of a. (a) 3 (b) 6 (c) 21 (d) 42 (e) the answer cannot be determined from the information given.

OpenStudy (lgbasallote):

two equal signs? is this written right?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

well i assume it means 3a + 4b = 4a - 4b subtract 3a from both sides 4b = a - 4b add 4b to both sides 8b = a = 21 so this is equal to 21? a = 21 8b = 21 b = 21/8 does that make sense? o.O

OpenStudy (anonymous):

i think you're supposed to add the 2 equations. then you get a is 6 and then b is 3/4

OpenStudy (anonymous):

a=6

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

and b=3/4

OpenStudy (lgbasallote):

like i said this question doesnt make sense -_-

OpenStudy (lgbasallote):

so you did 3a + 4b = 21 4a - 4b = 21 7a = 42 a = 6 ??

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

yeah that makes more sense...

OpenStudy (anonymous):

if u have doubts see here http://www.wolframalpha.com/input/?i=+3a+%2B+4b+%3D+4a+-+4b+%3D+21+find+a

OpenStudy (lgbasallote):

but seriously...two equal signs dont make sense :|

OpenStudy (anonymous):

it don't it would seem like the answer is E

OpenStudy (anonymous):

i mean (b) oops

OpenStudy (lgbasallote):

well if the question meant what you did...then 6 is right

OpenStudy (lgbasallote):

and also if you check 3(6) + 4(3/4) = 21 18 + 3 = 21 21 = 21 4(6) - 4(3/4) = 21 24 - 3 = 21 21 = 21

OpenStudy (lgbasallote):

so yeah. 6 would be right

OpenStudy (lgbasallote):

but still poor sense of notation =_=

OpenStudy (anonymous):

idk it was an sat question